# On the numerical stability of sketched GMRES

### v2

@@ Abstract
We perform a backward stability analysis of preconditioned sketched GMRES [Nakatsukasa and Tropp, SIAM J. Matrix Anal. Appl, 2024] for solving linear systems $Ax=b$, and show that the backward stability at iteration $i$ depends on the conditioning of the Krylov basis $B_{1:i}$ as long as the condition number of $AB_{1:i}$ can be bounded by $1/O(\bm{u})$, where $\bm{u}$ is the unit roundoff. Under this condition, we demonstrate that the stability of the sketched GMRES is mainly affected by the condition number of $B_{1:i}$, which can be mitigated by using the restarting technique. We also derive sharper bounds that explain why the backward error can be small even in cases when the basis $B_{1:i}$ is very ill-conditioned, which has been observed in the literature but not yet explained theoretically. We present numerical experiments to demonstrate the conclusions of our analysis, and also show that adaptively restarting where appropriate allows us to recover backward stability in sketched GMRES.
@@ 1 Introduction
In this paper, we analyze the numerical stability of sketched GMRES (sGMRES) [NT2024] for solving nonsymmetric and nonsingular linear systems of the form
@@
Although standard GMRES [SaadSchultz1986] is one of the most popular methods to solve (1), it comes at the expense of constructing a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace
@@
Sketched GMRES was proposed in [NT2024] as an attractive alternative to GMRES that exploits the technique of randomized sketching [martinsson2020randomized] to avoid the high cost of constructing an orthonormal basis. Instead, the basis is only partially orthogonalized, thereby reducing the arithmetic cost from $O(ni^{2})$ to $O(i^{3}+ni\log i)$.
It has been observed in numerical experiments in [NT2024] that sGMRES can fail once the condition number of the basis $B_{1:i}$ becomes too large, and methods such as the sketch-and-select Arnoldi algorithm [guttel2024sketch] have been proposed as a means to ensure the condition number of the basis $B_{1:i}$ does not grow too rapidly. Indeed, numerical experiments in [guttel2024sketch] showed that working with a well-conditioned basis improved the overall convergence behavior of sketched GMRES.
In this paper, we rigorously analyze the backward stability of preconditioned sketched GMRES, and justify the experimental observations made in the literature [NT2024, guttel2024sketch]. Namely, we show that the backward error of sketched GMRES depends on $\kappa(Z_{1:i})$ as long as $\kappa(AZ_{1:i})$ is not too large, where $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with the right preconditioner $M_{R}$. We then provide a backward stability result for a restarted version of sketched GMRES, which under certain assumptions, allows us to recover the numerical stability of sketched GMRES when $\kappa(Z_{1:i})$ grows large, suggesting we should monitor the conditioning of the basis throughout the iterations and restart when necessary. We present numerical experiments to demonstrate the numerical behavior we expect to observe in sketched GMRES as predicted by our new analysis, and demonstrate the improvements in the backward error that can be obtained from the adaptive restart approach.
@@
The remainder of this paper is organized as follows. In Section 2, we give an overview of GMRES and the MOD-GMRES framework [BHMV2024] for performing a backward stability analysis of different GMRES variants. We then discuss preconditioned sketched GMRES in Section 3, and analyze the backward stability of preconditioned sketched GMRES and restarted sketched GMRES, respectively, in Sections 4 and 5. Numerical experiments are presented in Section 6.
@@ 2 Preconditioned GMRES
There are many algorithmic variants of GMRES arising from different implementation options, such as the choice of preconditioner (if any), or the choice of orthogonalization method used within the Arnoldi procedure. Furthermore, for optimal performance it is common to exploit techniques such as deflation, mixed precision, or randomization within GMRES. In order to understand the finite precision behavior of any given GMRES variant, a backward error analysis framework was developed in [BHMV2024], which simplifies the process of deriving bounds for the attainable normwise backward and forward errors of the computed solutions of different GMRES variants. The framework is built from a backward error analysis of a generic GMRES algorithm known as modular GMRES or MOD-GMRES, that is composed of three elemental operations:
@@
Without loss of generality, we assume $x^{(0)}=0$ in this framework and our analysis of sGMRES in Section 4. By specializing these operations and meeting their assumptions, as outlined in the framework, modular GMRES can describe many GMRES implementations and variants, and can be used to perform their backward error analysis. For example, well-known results from [drkovsova1995numerical] were derived in [BHMV2024] using the framework. In particular, it was shown that GMRES with Householder orthogonalization (HH-GMRES) produces a computed solution whose backward error is on the order of the unit roundoff, and is thus backward stable. Similarly, the backward stability result for GMRES with modified Gram-Schmidt (MGS-GMRES), derived in [paige2006modified], was also recovered using the framework in [BHMV2024].
@@ 3 Preconditioned Sketched GMRES
for a preconditioned basis $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with a (non-orthogonal) Krylov basis $B_{1:i}$ and sketching operator $S\in\mathbb{R}^{s\times n}$ with $i<s\ll n$. The sketched problem (4) can then be constructed and solved inexpensively, even without orthonormality of the basis $B_{1:i}$. The solution of the sketched problem serves as a good approximation of the original problem as long as the matrix $S$ is chosen to be an $\epsilon$-subspace embedding for the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$, i.e., $S$ satisfies
$(1-\epsilon)\|v\|^{2}\leq\|Sv\|^{2}\leq(1+\epsilon)\|v\|^{2},$ (5)
for $\epsilon\in(0,1)$, and every $v\in\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$. In practice, the sketching matrix $S$ may not be explicitly available and needs to be drawn at random to achieve (5) with high probability. Additionally, $S$ must be constructed without prior knowledge of the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$ and should typically be a valid embedding for a broad subspace, specifically $\mathbb{R}^{n\times i}$. In practice, the sketched least squares problem (4) is computed by solving a triangular linear system
@@
It is clear from experimental observations made in [NT2024, guttel2024sketch] that the basis condition number should not grow too large throughout the iteration in order to ensure numerical stability of sketched GMRES. The non-orthogonal basis can be constructed using the $t$ truncated Arnoldi process, as summarized in Algorithm 3, where each new Krylov vector is orthogonalized against the previous $t$ vectors. Alternatively, the sketch-and-select method [guttel2024sketch] uses the sketched basis $SB_{i}$ to select $t$ vectors, among all previously computed vectors, with which to perform the partial orthogonalization, with the aim of generating a basis with smaller condition number than that generated by truncated Arnoldi.
@@ 4 Backward stability of preconditioned sketched GMRES
In this section, we aim to analyze the backward stability of preconditioned sketched GMRES (Algorithm 1) by primarily following the analysis of the MOD-GMRES framework [BHMV2024]. The first $i$ iterations of preconditioned sketched GMRES can be summarized in the following steps:
@@
with $\varepsilon_{W}\leq O(\bm{u})$ from the standard rounding error analysis of [H2002]. Similarly, we assume that the computed result of $\tilde{b}:=M_{L}^{-1}b$ computed by Line 1 in Algorithm 1 satisfies
@@
In iteratively solving the least squares problem in Step 2, we assume that there is a special dimension $k\leq n$, known as the key dimension [BHMV2024], for which the computed solution is guaranteed to have reached a small backward error, i.e., $\hat{b}$ lies in the range of non-singular basis $\hat{W}_{1:k}$. The definition of the key dimension is as follows:
@@
with $\varepsilon_{sls}\leq O(\bm{u})$ from [H2002, Theorem 20.3]. Substituting $\hat{C}_{1:i}$ and $\hat{g}$ in (13) with (11) and (12), we have
@@
We follow the proof in [BHMV2024] and employ [paige2006modified, Theorem 2.4] to bound $\|r^{(k)}_{S}\rVert$. Applying [paige2006modified, Theorem 2.4] to the least squares problem (14), we have
@@
We choose $\phi$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ as in [BHMV2024, (3.17)], which can simplify (19) as
@@
with $\varepsilon_{x}\leq O(\bm{u})$ from [H2002].
@@
Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:i}\in\mathbb{R}^{n\times i}$ and an $\epsilon$-subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (5) with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied. Then there exists $k\leq n$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$. If it further holds that
$\sqrt{k}\,c_{2}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq 1/2,\qquad\varepsilon_{x}\kappa(\hat{Z}_{1:k})\leq 1/2,$ (28)
@@
Note that $\hat{b}\in\mathbb{R}^{n}$, and we can establish the existence of $k$, since the condition $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$ is satisfied when $k=n$.
Then we only need to prove (29). The residual of $\hat{x}^{(k)}$ can be bounded by the preconditioned residual as follows:
@@
Furthermore, from [NT2024, Fact 2.2] and the definition (15) of $r^{(k)}_{S}$, we have
@@
It remains to bound $\|r^{(k)}_{S}\rVert$ and $\|\hat{y}^{(k)}\rVert$. To apply Lemma 1 to bound $\|r^{(k)}_{S}\rVert$, we first need to verify (16) and (17). From [BG2022, Corollary 2.2], the condition numbers of the sketched matrices satisfy
@@
Theorem 1 indicates that (29) and (30) are satisfied with high probability. This is due to the $\epsilon$-subspace embedding, which generally meets the condition (5) with high probability. Practically, the sketching method rarely fails, and slightly increasing the embedding dimension can further improve its reliability. In the worst case, $\epsilon$ may approach $1$ and the constant $(1+\epsilon)/(1-\epsilon)$ may be relatively large. For this reason, it is distracting to quantify the failure probabilities; see details in [MT2020].
@@
from (7). Therefore, if $S$ is chosen to be a $n\times n$ identity matrix and $\hat{B}_{1:k}$ is chosen to be an orthonormal Krylov basis, Theorem 1 recovers a similar backward stability result as [BHMV2024, Theorem 3.1].
@@
It has been observed in numerical experiments in [guttel2024sketch, Figure 9] that in some cases, sGMRES will continue to converge even when $\kappa(\hat{Z}_{1:k})>10^{15}$, and so the bound shown in Theorem 1 is not tight. Therefore, we present the following lemma to show a tighter bound, which can be directly derived from (37) by noticing
@@
Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:k}\in\mathbb{R}^{n\times k}$ and an $\epsilon$-subspace embedding $S\in\mathbb{R}^{s\times n}$, where (7), (8), (11), (12), (13), and (27) are satisfied. Then there exists $k\leq n$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$. If it further holds that
$\sqrt{k}\,c_{2}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq 1/2,$
@@
5 Adaptive restarting sGMRES

### v3

@@ Abstract
We perform a backward stability analysis of preconditioned sketched GMRES [Nakatsukasa and Tropp, SIAM J. Matrix Anal. Appl, 2024] for solving linear systems $Ax=b$, and show that the backward stability at iteration $i$ depends on the conditioning of the Krylov basis $B_{1:i}$ as long as the condition number of $AB_{1:i}$ can be bounded by $1/O(\bm{u})$, where $\bm{u}$ is the unit roundoff. Under this condition, we demonstrate that the stability of the sketched GMRES is mainly affected by the condition number of $B_{1:i}$, which can be mitigated by using the restarting technique. We also derive sharper bounds that better capture the attainable backward error especially for cases when the basis $B_{1:i}$ is very ill-conditioned, which has been observed in the literature but not yet explained theoretically. We present numerical experiments to demonstrate the conclusions of our analysis, and also show that adaptively restarting where appropriate allows us to recover backward stability in sketched GMRES.
@@ 1 Introduction
In this paper, we analyze the numerical stability of sketched GMRES (sGMRES) [10] for solving nonsymmetric and nonsingular linear systems of the form
@@
Although standard GMRES [12] is one of the most popular methods to solve (1), it comes at the expense of constructing a matrix $V_{1:i}\in\mathbb{R}^{n\times i}$ with columns that form an orthonormal basis for the Krylov subspace
@@
Sketched GMRES was proposed in [10] as an attractive alternative to GMRES that exploits the technique of randomized sketching [8] to avoid the high cost of constructing an orthonormal basis. Instead, the basis is only partially orthogonalized, thereby reducing the arithmetic cost from $O(ni^{2})$ to $O(i^{3}+ni\log i)$.
It has been observed in numerical experiments in [10] that sGMRES can fail once the condition number of the basis $B_{1:i}$ becomes too large, and methods such as the sketch-and-select Arnoldi algorithm [6] have been proposed as a means to ensure the condition number of the basis $B_{1:i}$ does not grow too rapidly. Indeed, numerical experiments in [6] showed that working with a well-conditioned basis improved the overall convergence behavior of sketched GMRES.
In this paper, we rigorously analyze the backward stability of preconditioned sketched GMRES, and justify the experimental observations made in the literature [10, 6]. Namely, we show that the backward error of sketched GMRES depends on $\kappa(Z_{1:i})$ as long as $\kappa(AZ_{1:i})$ is not too large, where $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with the right preconditioner $M_{R}$. We then provide a backward stability result for a restarted version of sketched GMRES, which under certain assumptions, allows us to recover the numerical stability of sketched GMRES when $\kappa(Z_{1:i})$ grows large, suggesting we should monitor the conditioning of the basis throughout the iterations and restart when necessary. We present numerical experiments to demonstrate the numerical behavior we expect to observe in sketched GMRES as predicted by our new analysis, and demonstrate the improvements in the backward error that can be obtained from the adaptive restart approach.
@@
The remainder of this paper is organized as follows. In Section 2, we give an overview of GMRES and the MOD-GMRES framework [3] for performing a backward stability analysis of different GMRES variants. We then discuss preconditioned sketched GMRES in Section 3, and analyze the backward stability of preconditioned sketched GMRES and restarted sketched GMRES, respectively, in Sections 4 and 5. Numerical experiments are presented in Section 6.
@@ 2 Preconditioned GMRES
There are many algorithmic variants of GMRES arising from different implementation options, such as the choice of preconditioner (if any), or the choice of orthogonalization method used within the Arnoldi procedure. Furthermore, for optimal performance it is common to exploit techniques such as deflation, mixed precision, or randomization within GMRES. In order to understand the finite precision behavior of any given GMRES variant, a backward error analysis framework was developed in [3], which simplifies the process of deriving bounds for the attainable normwise backward and forward errors of the computed solutions of different GMRES variants. The framework is built from a backward error analysis of a generic GMRES algorithm known as modular GMRES or MOD-GMRES, that is composed of three elemental operations:
@@
Without loss of generality, we assume $x^{(0)}=0$ in this framework and our analysis of sGMRES in Section 4. By specializing these operations and meeting their assumptions, as outlined in the framework, modular GMRES can describe many GMRES implementations and variants, and can be used to perform their backward error analysis. For example, well-known results from [5] were derived in [3] using the framework. In particular, it was shown that GMRES with Householder orthogonalization (HH-GMRES) produces a computed solution whose backward error is on the order of the unit roundoff, and is thus backward stable. Similarly, the backward stability result for GMRES with modified Gram-Schmidt (MGS-GMRES), derived in [11], was also recovered using the framework in [3].
@@ 3 Preconditioned Sketched GMRES
Before introducing sketched GMRES, we first recall the basic definition of the random sketching technique. For $\epsilon\in(0,1)$, a sketching matrix $S\in\mathbb{R}^{s\times n}$ is an $\epsilon$-subspace embedding for the subspace $\mathcal{V}\subseteq\mathbb{R}^{n\times m}$ if $S$ satisfies that
$(1-\epsilon)\|v\|^{2}\leq\|Sv\|^{2}\leq(1+\epsilon)\|v\|^{2},$ (4)
for any $v\in\mathcal{V}$. In practice, the sketching matrix $S$ may not be explicitly available and needs to be drawn at random to achieve (4) with high probability.
@@
for a preconditioned basis $Z_{1:i}=M_{R}^{-1}B_{1:i}$ with a (non-orthogonal) Krylov basis $B_{1:i}$ and sketching operator $S\in\mathbb{R}^{s\times n}$ with $i<s\ll n$. The sketched problem (5) can then be constructed and solved inexpensively, even without orthonormality of the basis $B_{1:i}$. The solution of the sketched problem serves as a good approximation of the original problem as long as the matrix $S$ is chosen to be an $\epsilon$-subspace embedding for the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$, i.e., $S$ satisfies (4) for $\epsilon\in(0,1)$, and every $v\in\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$. Additionally, $S$ must be constructed without prior knowledge of the Krylov subspace $\mathcal{K}_{i}(M_{L}^{-1}AM_{R}^{-1},M_{L}^{-1}r^{(0)})$ and should typically be a valid embedding for a broad subspace, specifically $\mathbb{R}^{n\times i}$. In practice, the sketched least squares problem (5) is computed by solving a triangular linear system
@@
It is clear from experimental observations made in [10, 6] that the basis condition number should not grow too large throughout the iteration in order to ensure numerical stability of sketched GMRES. The non-orthogonal basis can be constructed using the $t$ truncated Arnoldi process, as summarized in Algorithm 3, where each new Krylov vector is orthogonalized against the previous $t$ vectors. Alternatively, the sketch-and-select method [6] uses the sketched basis $SB_{i}$ to select $t$ vectors, among all previously computed vectors, with which to perform the partial orthogonalization, with the aim of generating a basis with smaller condition number than that generated by truncated Arnoldi.
@@ 4 Backward stability of preconditioned sketched GMRES
In this section, we aim to analyze the backward stability of preconditioned sketched GMRES (Algorithm 1) by primarily following the analysis of the MOD-GMRES framework [3]. The first $i$ iterations of preconditioned sketched GMRES can be summarized in the following steps:
@@
with $\varepsilon_{W}\leq O(\bm{u})$ from the standard rounding error analysis of [7]. Similarly, we assume that the computed result of $\tilde{b}:=M_{L}^{-1}b$ computed by Line 1 in Algorithm 1 satisfies
@@
In iteratively solving the least squares problem in Step 2, we assume that there is a special dimension $k\leq n$, known as the key dimension [3], for which the computed solution is guaranteed to have reached a small backward error, i.e., $\hat{b}$ lies in the range of non-singular basis $\hat{W}_{1:k}$. The definition of the key dimension is as follows:
@@
with $\varepsilon_{sls}\leq O(\bm{u})$ from [7, Theorem 20.3]. Substituting $\hat{C}_{1:i}$ and $\hat{g}$ in (13) with (11) and (12), we have
@@
We follow the proof in [3] and employ [11, Theorem 2.4] to bound $\|r^{(k)}_{S}\rVert$. Applying [11, Theorem 2.4] to the least squares problem (14), we have
@@
We choose $\phi$ such that $\phi^{-2}=\|\hat{y}^{(k)}\rVert^{2}/(1-\delta_{k}^{2})$ as in [3, (3.17)], which can simplify (19) as
@@
with $\varepsilon_{x}\leq O(\bm{u})$ from [7].
@@
Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:m}\in\mathbb{R}^{n\times m}$ and a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (4) for the subspace $\mathbb{R}^{n\times m}$ with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied. If there exists $k\leq m$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$,
$\sqrt{k}\,c_{2}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq\frac{1}{2},\quad\text{and}\quad\varepsilon_{x}\kappa(\hat{Z}_{1:k})\leq\frac{1}{2},$ (28)
@@
The residual of $\hat{x}^{(k)}$ can be bounded by the preconditioned residual as follows:
@@
Furthermore, from [10, Fact 2.2] and the definition (15) of $r^{(k)}_{S}$, we have
@@
It remains to bound $\|r^{(k)}_{S}\rVert$ and $\|\hat{y}^{(k)}\rVert$. To apply Lemma 1 to bound $\|r^{(k)}_{S}\rVert$, we first need to verify (16) and (17). From [2, Corollary 2.2], the condition numbers of the sketched matrices satisfy
@@
Theorem 1 indicates that (29) and (30) are satisfied with high probability. This is due to the choice of the subspace embedding $S$, which generally meets the condition (4) with high probability. Practically, the sketching method rarely fails, and slightly increasing the embedding dimension can further improve its reliability. In the worst case, $\epsilon$ may approach $1$ and the constant $(1+\epsilon)/(1-\epsilon)$ may be relatively large. For this reason, it is distracting to quantify the failure probabilities; see details in [9].
Remark 2.
For standard GMRES, the existence of such a $k$ is straightforward to be guaranteed by observing that $\hat{b}\in\mathbb{R}^{n}$, because the condition $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\,\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$ is automatically satisfied when $k=n$. In contrast, for sketched GMRES, one must select a proper subspace embedding $S\in\mathbb{R}^{s\times n}$ for the $m$-dimensional subspace in order to ensure the existence of such a $k$. When $s$ is sufficiently large—particularly in the extreme case s=ns=n—the existence of $k$ can be guaranteed in essentially the same way as for standard GMRES, but then the advantage of employing random sketching is lost.
@@
from (7). Therefore, if $S$ is chosen to be a $n\times n$ identity matrix and $\hat{B}_{1:k}$ is chosen to be an orthonormal Krylov basis, Theorem 1 recovers a similar backward stability result as [3, Theorem 3.1].
@@
It has been observed in numerical experiments in [6, Figure 9] that in some cases, sGMRES will continue to converge even when $\kappa(\hat{Z}_{1:k})>10^{15}$, and so the bound shown in Theorem 1 is not tight. Therefore, we present the following lemma to show a tighter bound, which can be directly derived from (37) by noticing
@@
Assume that Algorithm 1 is applied with a computed preconditioned basis $\hat{Z}_{1:m}\in\mathbb{R}^{n\times m}$ and a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (4) for the subspace $\mathbb{R}^{n\times m}$ with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied. If there exists $k\leq m$ such that $c_{1}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa(\begin{bmatrix}\phi\hat{b}&\hat{W}_{1:k}\end{bmatrix})\geq 1$,
$\sqrt{k}\,c_{2}(n,k)\bigl(\varepsilon_{S}+\varepsilon_{sls}\bigr)\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k})\leq\frac{1}{2},\quad\text{and}\quad\varepsilon_{x}\kappa(\hat{Z}_{1:k})\leq\frac{1}{2},$
@@
5 Restarted sGMRES with adaptive truncation size